OpenStudy (anonymous):

What is the product of (n+6)(n-6)

OpenStudy (anonymous):

n^2-36

OpenStudy (anonymous):

depends what n is , but it is n^2-36 if you ask me

OpenStudy (diyadiya):

it is of the form (a+b)(a-b) so \[(a+b)(a-b)=a^2-b^2\] \[(n+6)(n-6)=n^2-6^2=n^2-36\]